### v3

@@ 3.2 Implications to Stochastic Weakly Convex Minimizations
In the above, $\xi_{1},\xi_{2},\ldots,\xi_{m}$ are i.i.d. samples drawn from the distribution $\mathbb{P}$, where $\mathbb{P}_{m}$ denotes the empirical distribution of $\xi_{1},\xi_{2},\ldots,\xi_{m}$. Formally, we start with a single random variable $\xi$ defined on a probability space $(\Xi,\mathcal{G},\mathbb{P})$. Utilizing this space, we can construct the countable product space $(\Xi,\mathcal{G},\mathbb{P})^{\mathbb{N}}$ where the sequence of i.i.d. random variables $(\xi_{1},\xi_{2},\ldots,\xi_{m},\ldots)$ can be properly defined on this product probability space [30, Section 36], with each $\xi_{i}$ acting as a coordinate projection. In this product space, we then extend the sigma-algebra and the probability measure to obtain a complete measure, following standard procedures in measure theory [30], to avoid some cumbersome measurability issues [2, Section 7]. The function $R:\mathbb{R}^{d}\to\mathbb{R}$ is a real-valued, closed and convex function. The constraint set $X$ is a nonempty, closed, and convex set.
@@
is convex on the open ball $\mathbb{B}(x;\epsilon(x))$. The collection of these balls, $\{\mathbb{B}(x;\epsilon(x))\}_{x\in\mathcal{O}}$ forms an open cover of $\mathcal{O}$. By Lindelöf’s covering Theorem, see, e.g., [35, Page 50], and noting that $\mathbb{R}^{d}$ with its norm topology is second countable, there is a countable subcover $\{\mathbb{B}(x;\epsilon(x))\}_{x\in\mathcal{O}^{o}}$ of $\mathcal{O}$, where $\mathcal{O}^{o}$ is a countable subset of $\mathcal{O}$. In particular, $\mathcal{O}$ obeys the inclusion:
@@
3.3 A Class of Stochastic Weakly Convex Minimizations
One important example of stochastic weakly convex minimizations corresponds to the setting where
@@ 3.3 A Class of Stochastic Weakly Convex Minimizations
Here, $h:\mathbb{R}^{k}\to\mathbb{R}$ is a closed, convex and real-valued function, and $c(\cdot,\xi):\mathbb{R}^{d}\to\mathbb{R}^{k}$ is $\mathcal{C}^{1}$ smooth on $\mathbb{R}^{d}$ for every $\xi$. Under this scenario, the population and empirical objectives become
@@
We give examples of stochastic convex-composite objectives that appear in statistics, machine learning, imaging, and risk management. More examples can be found in [36, Section 2.1].
@@
Phase retrieval is a computational problem with applications across various fields, including imaging, X-ray crystallography, and speech processing. The (real-valued) phase retrieval seeks to detect a point $x$ satisfying $|\langle a_{i},x\rangle|^{2}=b_{i}$ where $a_{i}\in\mathbb{R}^{d}$, and $b_{i}\in\mathbb{R}$ for $i=1,2,\ldots,m$. We can choose $\xi=(a,b)\in\mathbb{R}^{d}\times\mathbb{R}$, $h(z)=|z|$ and $c(x;\xi)=(a^{T}x)-b$, and $X=\mathbb{R}^{d}$, in which case the form (3.9) gives an exact penalty formulation for solving the collection of quadratic equations, which yields strong statistical recovery and robustness guarantees [37, 8, 14], among other nonconvex formulations [38, 39, 40].
@@
  • $\sup_{y\in\mathbb{B}(x;\epsilon(x))}\left\|{\nabla c(y;\xi)\partial h(c(y;\xi))}\right\|$ is integrable with respect to $\xi\sim\mathbb{P}$.
@@
For ease of reference in future discussions, we establish a result that stems directly from Theorem 2 and Lemma 2 for stochastic convex-composite minimizations.
@@ 4 Explicit Uniform Convergence Rates
Here, the crucial assumption is that $h:\mathbb{R}\to\mathbb{R}$ is a one-dimensional convex function. The function $c(\cdot;\xi):\mathbb{R}^{d}\to\mathbb{R}$ is $\mathcal{C}^{1}$ smooth on $\mathbb{R}^{d}$ for every $\xi$, aligning with the framework in Section 3.3. As before, $R:\mathbb{R}^{d}\to\mathbb{R}$ is closed and convex, and $X$ is nonempty, closed and convex. This setting is particularly relevant as it covers a broad spectrum of practical objectives, including all examples previously discussed in Section 3.3. In our considerations, the set $\mathcal{O}$ is usually an open ball $\mathbb{B}(x_{0};r)$ in the Euclidean space where $x_{0}\in\mathbb{R}^{d}$ is the center and $r>0$ is the radius.
@@ 4.1 Subgradient Selections
This selection process begins with identifying a subgradient for the nonsmooth convex function $h$, which we decompose into a sum $h=h^{sm}+h^{ns}$ of a smooth component $h^{sm}$ and a nonsmooth component $h^{ns}$. To do so, we use $h^{\prime}_{+}$ and $h^{\prime}_{-}$ to denote the right-hand and left-hand derivative of $h$, respectively. We enumerate the set of nondifferentiable points $\{t\in\mathbb{R}:h^{\prime}_{+}(t)\neq h^{\prime}_{-}(t)\}$ as $\{t_{j}\}_{j=1}^{\infty}$, given the fact that these nondifferentiable points are always countable [42, Theorem 2.1.2].
@@
$h(x)=h^{sm}(z)+h^{ns}(z).$ (4.5)
Here, $h^{ns}:\mathbb{R}\mapsto\mathbb{R}$ (where “ns” stands for nonsmooth) denotes the nonsmooth component of $h$, which is defined by:
$h^{ns}(x)=\sum_{j=1}^{\infty}a_{j}\cdot(z-t_{j})_{+},$ (4.6)
where the coefficients are defined by $a_{j}=h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j})>0$. Notably, $h^{ns}$ is well-defined because $\sum_{j=1}^{\infty}a_{j}=\sum_{j=1}^{\infty}(h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j}))<\infty$. The smooth component $h^{sm}:\mathbb{R}\mapsto\mathbb{R}$ (where “sm” stands for smooth) is defined as $h^{sm}=h-h^{ns}$. By construction, $h^{sm}$ is a $\mathcal{C}^{1}$ smooth and convex function, as one can easily verify its derivative $(h^{sm})^{\prime}$ is continuous and monotonically increasing on $\mathbb{R}$.
@@
To summarize, $G$ and $G_{S}$ are well-defined whenever $\sum_{j=1}^{\infty}(h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j}))<\infty$.
@@ 4.2 A Uniform Bound on Subgradient Selections
We state the definition of VC dimension for completeness [22, Section 3.6].
@@
=$\mathcal{F}=\{\{\xi\in\Xi:c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}.$ (4.12)
@@
The complexity measure ${\rm vc}(\mathcal{F})$ characterizes generalization capabilities of $\mathcal{F}$ and bounds the uniform rate of convergence from $G_{S}$ to $G$. For common techniques of upper bounding VC dimensions of set families, see [31, Chapter 2]. When the mapping $x\mapsto c(x;\xi)$ are polynomials in $x\in\mathbb{R}^{d}$ for any given $\xi$—a condition met in our applications—we establish a result giving a tight upper bound of the VC dimension, based on results in real algebraic geometry. See Section 6 for details.
Our main results rely on Lipschitz continuity and regularity assumptions concerning the tails of the random vectors $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$. These random vectors appear in our subgradient selection (4.11), and our assumptions about them align with standard treatments in the literature. To state the assumptions, we first recall the concept of a subexponential tail random vector. We define this in terms of Orlicz norms, following [44, Section 3.4.4].
@@
Examples of $\sigma$-sub-exponential random vectors include a normal distribution $Z\sim\mathsf{N}(0,\Sigma)$ where the largest eigenvalue of $\Sigma$ is bounded above by $c\sigma$, or distributions with bounded support, such as $\left\|{Z}\right\|\leq c\sigma$ almost surely, with the numerical constant $c=1/2$ [44, Section 2].
By Markov’s inequality, $\mathbb{P}(|\langle v,Z\rangle|\geq\sigma\cdot u)\leq 2e^{-u}$ for any unit vector $v$ and $u\geq 0$. Furthermore, by Bernstein’s inequality [44, Theorem 2.8.1] and the centering property of sub-exponential random variables [44, Exercise 2.7.10], for i.i.d. $\sigma$-subexponential random variables $Z_{1},Z_{2},\ldots,Z_{m}$, there is the exponential tail bound for the mean $\frac{1}{m}\sum_{i=1}^{m}Z_{i}$ that holds for every $v$ with $\left\|{v}\right\|=1$, and $u\geq 0$:
@@
This assumption is commonly used in empirical process theory to extend high probability bounds from individual points to the uniform control of the supremum of stochastic processes, utilizing the chaining technique. Notably, this assumption naturally holds in a variety of statistical applications. For a reference, see [31, Chapter 2].
We are now ready to present our main result. The proof of Theorem 4 employs empirical process theory tools, namely, the calculus rules for outer probability measure [31, Chapter 1.2-5], the chaining argument [45], and the Sauer-Shelah Lemma—a fundamental combinatorial principle using the VC dimension to control function class complexities [46, 47]. The proof is relatively standard for experts in statistical learning theory and is detailed in Section A.
@@
Finally, the $\log m$ term arises from the subexponential tail assumption and can be removed if a norm boundedness assumption replaces it in Assumption C.2, or by using more advanced arguments involving additional probabilistic control of the envelope function of $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$ [31, Chapter 2]. For conciseness and clarity, this paper does not pursue such removal.
@@ 4.3 Implication: A Uniform Bound on Subdifferentials
We integrate Theorem 3 and Theorem 4 to document our ultimate result, Theorem 5, on uniform convergence of subdifferentials. At this point, the only work is to collect and streamline all the assumptions in Theorem 3 and 4. To fulfill Assumption B’ required in Theorem 3, we identify an additional condition beyond Assumptions C.1C.2. This Lipschitz-type condition is very mild, and holds in many applications in stochastic programming [2, Chapter 5] and statistics [49, Chapter 5].
@@
hold for every $x_{1},x_{2}\in\mathbb{B}(x_{0};r)$, where $e(x,\xi)=(\psi^{sm})^{\prime}(c(x,\xi))\nabla c(x,\xi)$.
@@
Assumptions C.1C.2 imply that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)}\right\|$ and $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{(h^{sm})^{\prime}(c(y;\xi))\nabla c(y;\xi)}\right\|$ are integrable over $\xi\sim\mathbb{P}$ due to the standard chaining argument [44, Section 8]. This implies that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)\partial h(c(y;\xi))}\right\|$ is also integrable over $\xi$ since $\sum_{j=1}^{\infty}(h_{+}^{\prime}(t_{j})-h_{-}^{\prime}(t_{j}))<\infty$, thus confirming the first condition of Assumption B’ holds for $\mathcal{O}=\mathbb{B}(x_{0};r)$. The second condition of Assumption B’ follows trivially from Assumption C.3. ∎
@@
Theorem 5 establishes a rate for uniform convergence of subdifferential mappings for stochastic convex-composite objectives at $\sqrt{\max\{d,{\rm vc}(\mathcal{F})\}/m}$, modulo logarithmic factors. Our results are different from approaches that depend on the continuity of both empirical and population subdifferentials under the Hausdorff metric, or require a nonatomic probability distribution $\mathbb{P}$ to achieve a $1/\sqrt{m}$ rate [1, 21] (see Section 1.1 for a more detailed literature review). Indeed, Theorem 5 achieves this tight rate without requiring continuity in the subdifferential mappings of the objective or specific conditions on the data distribution $\mathbb{P}$.

### v4

@@ 3.2 Implications to Stochastic Weakly Convex Minimization Problems
In the above, $\xi_{1},\xi_{2},\ldots,\xi_{m}$ are i.i.d. samples drawn from the distribution $\mathbb{P}$, where $\mathbb{P}_{m}$ denotes the empirical distribution of $\xi_{1},\xi_{2},\ldots,\xi_{m}$. Formally, we start with a single random variable $\xi$ defined on a probability space $(\Xi,\mathcal{G},\mathbb{P})$. Utilizing this space, we can construct the countable product space $(\Xi,\mathcal{G},\mathbb{P})^{\mathbb{N}}$ where the sequence of i.i.d. random variables $(\xi_{1},\xi_{2},\ldots,\xi_{m},\ldots)$ can be properly defined on this product probability space [33, Section 36], with each $\xi_{i}$ acting as a coordinate projection. In this product space, we then extend the sigma-algebra and the probability measure to obtain a complete measure, following standard procedures in measure theory [33], to avoid some cumbersome measurability issues [2, Section 7]. The function $R\colon\mathbb{R}^{d}\to\mathbb{R}$ is a real-valued, closed and convex function. The constraint set $X$ is a nonempty, closed, and convex set.
@@
is convex on the open ball $\mathbb{B}(x;\epsilon(x))$. The collection of these balls, $\{\mathbb{B}(x;\epsilon(x))\}_{x\in\mathcal{O}}$ forms an open cover of $\mathcal{O}$. By Lindelöf’s covering Theorem, see, e.g., [36, Page 50], and noting that $\mathbb{R}^{d}$ with its norm topology is second countable, there is a countable subcover $\{\mathbb{B}(x;\epsilon(x))\}_{x\in\mathcal{O}^{o}}$ of $\mathcal{O}$, where $\mathcal{O}^{o}$ is a countable subset of $\mathcal{O}$. In particular, $\mathcal{O}$ obeys the inclusion:
@@
3.3 A Class of Stochastic Weakly Convex Minimization Problems
One important example of stochastic weakly convex minimization problems corresponds to the setting where
@@ 3.3 A Class of Stochastic Weakly Convex Minimization Problems
Here, $h\colon\mathbb{R}^{k}\to\mathbb{R}$ is a closed, convex and real-valued function, and $c(\cdot,\xi)\colon\mathbb{R}^{d}\to\mathbb{R}^{k}$ is $\mathcal{C}^{1}$ smooth on $\mathbb{R}^{d}$ for every $\xi$. Under this scenario, the population and empirical objectives become
@@
We give examples of stochastic convex-composite objectives that appear in statistics, machine learning, imaging, and risk management. More examples can be found in [37, Section 2.1].
@@
Phase retrieval is a computational problem with applications across various fields, including imaging, X-ray crystallography, and speech processing. The (real-valued) phase retrieval seeks to detect a point $x$ satisfying $|\langle a_{i},x\rangle|^{2}=b_{i}$ where $a_{i}\in\mathbb{R}^{d}$, and $b_{i}\in\mathbb{R}$ for $i=1,2,\ldots,m$. We can choose $\xi=(a,b)\in\mathbb{R}^{d}\times\mathbb{R}$, $h(z)=|z|$ and $c(x;\xi)=(a^{T}x)^{2}-b$, and $X=\mathbb{R}^{d}$, in which case the form (3.12) gives an exact penalty formulation for solving the collection of quadratic equations, which yields strong statistical recovery and robustness guarantees [38, 8, 14], among other nonconvex formulations [39, 40, 41].
@@
Example 3 (Robust Blind Deconvolution).
Blind deconvolution aims to recover two vectors $x_{1},x_{2}\in\mathbb{R}^{d}$ from the observed measurements $b_{i}=\langle u_{i},x_{1}\rangle\langle v_{i},x_{2}\rangle$ from $i=1,2,\ldots,n$ where $u_{i},v_{i}\in\mathbb{R}^{d}$ are known measurement vectors. This problem has applications in fields like astronomy and computer vision. A recent formulation for recovering $x_{1}$ and $x_{2}$ minimizes the reconstruction error [9, 43]:
$\min_{x_{1}\in\mathbb{R}^{d},x_{2}\in\mathbb{R}^{d}}\frac{1}{m}\sum_{i=1}^{m}|b_{i}-\langle u_{i},x_{1}\rangle\langle v_{i},x_{2}\rangle|$
This falls into the form (3.12) by simply setting $x=(x_{1},x_{2})$, $\xi=(u,v,b)$ where $u,v\in\mathbb{R}^{d},b\in\mathbb{R}$, $h(z)=|z|$ and $c(x;\xi)=c((x_{1},x_{2});(u,v,b))=b-\langle u,x_{1}\rangle\langle v,x_{2}\rangle$.
@@
  • $\sup_{y\in\mathbb{B}(x;\epsilon(x))}\left\|{\nabla c(y;\xi)\partial h(c(y;\xi))}\right\|\leq L_{x}(\xi)$ for some random variable $L_{x}(\xi)$ that is measurable, and integrable with respect to $\xi\sim\mathbb{P}$.
@@
For ease of reference in future discussions, we establish a result that stems directly from Theorem 2 and Lemma 2 for stochastic convex-composite minimization problems.
@@ 4 Explicit Uniform Convergence Rates
Here, the crucial assumption is that $h\colon\mathbb{R}\to\mathbb{R}$ is a one-dimensional convex function. The function $c(\cdot;\xi)\colon\mathbb{R}^{d}\to\mathbb{R}$ is $\mathcal{C}^{1}$ smooth on $\mathbb{R}^{d}$ for every $\xi$, aligning with the framework in Section 3.3. As before, $R\colon\mathbb{R}^{d}\to\mathbb{R}$ is closed and convex, and $X$ is nonempty, closed and convex. This setting is particularly relevant as it covers a broad spectrum of practical objectives, including all examples previously discussed in Section 3.3. In our considerations, the set $\mathcal{O}$ is usually an open ball $\mathbb{B}(x_{0};r)$ in the Euclidean space where $x_{0}\in\mathbb{R}^{d}$ is the center and $r>0$ is the radius.
@@ 4.1 Subgradient Selections
This selection process begins with identifying a subgradient for the nonsmooth convex function $h$, which we decompose into a sum $h=h^{sm}+h^{ns}$ of a smooth component $h^{sm}$ and a nonsmooth component $h^{ns}$. To do so, we use $h^{\prime}_{+}$ and $h^{\prime}_{-}$ to denote the right-hand and left-hand derivative of $h$, respectively. We enumerate the set of nondifferentiable points $\{t\in\mathbb{R}:h^{\prime}_{+}(t)\neq h^{\prime}_{-}(t)\}$ as $\{t_{j}\}_{j=1}^{\infty}$, given the fact that these nondifferentiable points are always countable [44, Theorem 2.1.2].
@@
$h(z)=h^{sm}(z)+h^{ns}(z).$ (4.5)
Here, $h^{ns}\colon\mathbb{R}\to\mathbb{R}$ (where “ns” stands for nonsmooth) denotes the nonsmooth component of $h$, which is defined by:
$h^{ns}(z)=\sum_{j=1}^{\infty}a_{j}\cdot(z-t_{j})_{+},$ (4.6)
where the coefficients are defined by $a_{j}=h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j})>0$. Notably, $h^{ns}$ is well-defined because $\sum_{j=1}^{\infty}a_{j}=\sum_{j=1}^{\infty}(h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j}))<\infty$. The smooth component $h^{sm}:\mathbb{R}\to\mathbb{R}$ (where “sm” stands for smooth) is defined as $h^{sm}=h-h^{ns}$. By construction, $h^{sm}$ is a $\mathcal{C}^{1}$ smooth and convex function, as one can easily verify its derivative $(h^{sm})^{\prime}$ is continuous and monotonically increasing on $\mathbb{R}$.
@@
To summarize, $G$ and $G_{S}$ are well-defined functions under the condition $\sum_{j=1}^{\infty}(h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j}))<\infty$, provided that regularity Assumptions A’ and B’ are also satisfied.
@@ 4.2 A Uniform Bound on Subgradient Selections
We state the definition of VC dimension for completeness [25, Section 3.6].
@@
=$\mathcal{F}=\{\{\xi\in\Xi\colon c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}.$ (4.12)
@@
The complexity measure ${\rm vc}(\mathcal{F})$ characterizes generalization capabilities of $\mathcal{F}$ and bounds the uniform rate of convergence from $G_{S}$ to $G$. For common techniques of upper bounding VC dimensions of set families, see [34, Chapter 2]. When the mapping $x\mapsto c(x;\xi)$ are polynomials in $x\in\mathbb{R}^{d}$ for any given $\xi$—a condition met in our applications—we establish a result giving a tight upper bound of the VC dimension, based on results in real algebraic geometry. See Section 5 for details.
Our main results rely on Lipschitz continuity and regularity assumptions concerning the tails of the random vectors $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$. These random vectors appear in our subgradient selection (4.11), and our assumptions about them align with standard treatments in the literature. To state the assumptions, we first recall the concept of a subexponential tail random vector. We define this in terms of Orlicz norms, following [46, Section 3.4.4].
@@
Examples of $\sigma$-sub-exponential random vectors include a normal distribution $Z\sim\mathsf{N}(0,\Sigma)$ where the largest eigenvalue of $\Sigma$ is bounded above by $c\sigma$, or distributions with bounded support, such as $\left\|{Z}\right\|\leq c\sigma$ almost surely, with the numerical constant $c=1/2$ [46, Section 2].
By Markov’s inequality, $\mathbb{P}(|\langle v,Z\rangle|\geq\sigma\cdot u)\leq 2e^{-u}$ for any unit vector $v$ and $u\geq 0$. Furthermore, by Bernstein’s inequality [46, Theorem 2.8.1] and the centering property of sub-exponential random variables [46, Exercise 2.7.10], for i.i.d. $\sigma$-subexponential random variables $Z_{1},Z_{2},\ldots,Z_{m}$, there is the exponential tail bound for the mean $\frac{1}{m}\sum_{i=1}^{m}Z_{i}$ that holds for every $v$ with $\left\|{v}\right\|=1$, and $u\geq 0$:
@@
This assumption is commonly used in empirical process theory to extend high probability bounds from individual points to the uniform control of the supremum of stochastic processes, utilizing the chaining technique. Notably, this assumption naturally holds in a variety of statistical applications. For a reference, see [34, Chapter 2].
We are now ready to present our main result. The proof of Theorem 4 employs empirical process theory tools, namely, the calculus rules for outer probability measure [34, Chapter 1.2-5], the chaining argument [47], and the Sauer-Shelah Lemma—a fundamental combinatorial principle using the VC dimension to control function class complexities [48, 49]. The proof is relatively standard for experts in statistical learning theory and is detailed in Section A.
@@
Finally, the $\log m$ term arises from the subexponential tail assumption and can be removed if a norm boundedness assumption replaces it in Assumption C.2, or by using more advanced arguments involving additional probabilistic control of the envelope function of $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$ [34, Chapter 2]. For conciseness and clarity, this paper does not pursue such removal.
@@ 4.3 Implication: A Uniform Bound on Subdifferentials
We integrate Theorem 3 and Theorem 4 to document our ultimate result, Theorem 5, on uniform convergence of subdifferentials. At this point, the only work is to collect and streamline all the assumptions in Theorem 3 and 4. To fulfill Assumption B’ required in Theorem 3, we identify an additional condition beyond Assumptions C.1C.2. This Lipschitz-type condition is very mild, and holds in many applications in stochastic programming [2, Chapter 5] and statistics [51, Chapter 5].
@@
hold for every $x_{1},x_{2}\in\mathbb{B}(x_{0};r)$, where $e(x,\xi)=(h^{sm})^{\prime}(c(x,\xi))\nabla c(x,\xi)$.
@@
Assumptions C.1C.2 imply that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)}\right\|$ and $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{(h^{sm})^{\prime}(c(y;\xi))\nabla c(y;\xi)}\right\|$ are integrable over $\xi\sim\mathbb{P}$ due to the standard chaining argument [46, Section 8]. This implies that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)\partial h(c(y;\xi))}\right\|$ has an integrable upper bound given that $\sum_{j=1}^{\infty}(h_{+}^{\prime}(t_{j})-h_{-}^{\prime}(t_{j}))<\infty$, thus confirming the first condition of Assumption B’ holds for $\mathcal{O}=\mathbb{B}(x_{0};r)$. The second condition of Assumption B’ follows trivially from Assumption C.3. ∎
@@
Theorem 5 establishes a rate for uniform convergence of subdifferential mappings for stochastic convex-composite objectives at $\sqrt{\max\{d,{\rm vc}(\mathcal{F})\}/m}$, modulo logarithmic factors. Our results are different from approaches that depend on the continuous differentiability of the population objectives, or require a nonatomic probability distribution $\mathbb{P}$ to achieve a $1/\sqrt{m}$ rate [1, 24] (see Section 1.1 for a more detailed literature review). Indeed, Theorem 5 achieves this tight rate without requiring continuity type assumptions in the subdifferential mappings of the objective or specific conditions on the data distribution $\mathbb{P}$.
@@
5 VC Dimension Bounds
In Theorem 6 below, we establish an upper bound on the VC dimension of the set $\mathcal{F}$ defined in equation (4.12) under the condition that the functions $x\mapsto c(x;\xi)$ are polynomials in $x$ for all $\xi$. Notably, our result does not require $\xi\mapsto c(x;\xi)$ to be polynomials in the data $\xi$. This general bound will be particularly useful for future applications of Theorem 5 to concrete problems such as matrix sensing, phase retrieval, and blind deconvolution, where $x\mapsto c(x;\xi)$ naturally takes the form of polynomials in $x$. For further details on these applications, see Section 6.
Theorem 6.
=$\mathcal{F}=\{\{\xi\in\Xi\colon c(x,\xi)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}$
is bounded by $Cd\log(Kd)$, where $C>0$ is a universal constant.
Definition 3.
The following theorem, rooted in the field of real algebraic geometry, is referenced from Matousek’s monograph on discrete geometry [52]. See [53, 54] and [55, Theorem 3].
Theorem 7 ([52, Theorem 6.2.1]).
$\left(\frac{50KN}{d}\right)^{d}.$
Proof of Theorem 6.
Consider the set of binary patterns $b$ defined by:
=$\begin{split}\mathcal{B}&=\{(b_{1},b_{2},\ldots,b_{N})\in\{0,1\}^{N}\colon\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{there exists $z\in\mathbb{R}^{d+1}$ such that}~{}b_{i}=\mathbf{1}\{q_{i}(z)\geq 0\}~{}\text{for all $1\leq i\leq N$}\}.\end{split}$
$|\mathcal{B}|\leq\left(\frac{50KN}{d+1}\right)^{d+1}.$
$\left(\frac{50KN}{d+1}\right)^{d+1}<2^{N}.$